# High-precision measurements of electric-dipole-transition amplitudes in excited states of 208Pb using Faraday rotation spectroscopy

### v1

@@ Abstract
We have completed measurements of two low-lying excited-state electric-dipole (E1) transition amplitudes in lead. Our measured reduced matrix elements of the $(6s^{2}6p^{2})^{3}P_{1}\to(6s^{2}6p7s)^{3}P_{0}$ transition at 368.3 nm and the 405.8 nm $(6s^{2}6p^{2})^{3}P_{2}\to(6s^{2}6p7s)^{3}P_{1}$ transition are 1.90(1) a.u. and 3.01(2) a.u. respectively, both measured to sub-1% precision and both in excellent agreement with the latest ab initio lead wavefunction calculations. These measurements were completed by comparing the low-field Faraday optical rotation spectra of each E1 transition in turn with that of the ground-state ${}^{3}P_{0}\to^{3}P_{1}$ M1 transition under identical experimental conditions. Our spectroscopy technique involves polarization modulation and lock-in detection yielding microradian-level optical rotation resolution. At temperatures where direct absorption was significant for both E1 and M1 transitions, we also extracted matrix element values from a direct optical absorption depth comparison. As part of this work we designed an interaction region within our furnace which allowed precise determination of our quartz vapor cell sample temperature to provide accurate determination of the Boltzmann thermal population of the low-lying excited states that were studied.
@@ I Introduction
Atoms and associated precision spectroscopic techniques have long been an excellent means to provide important, low-energy tests of fundamental physics associated with the standard model, as well as tests of potential new (“BSM”) physics beyond it DeMille2017 ; Safronova2018 . Heavy (high-$Z$, often multi-valence) atomic species in particular, due to their enhanced sensitivity to many of these phenomena, have played an important role in experimental searches. This high-precision experimental work has developed alongside ab initio atomic theory work which is required to link experimental results to interpretations and limits of the fundamental physics phenomena being targeted. Recent atomic theory work particularly targeted at multi-valence atomic systems, and broadly applicable to many such systems, has made notable progress in accuracy in recent years Safronova2009 ; Safronova2013 ; Safronova2025 . Unlike, for example, the alkali atom systems, far fewer general high-precision atomic structure measurements have been completed in heavy, multi-valence atoms with which to benchmark ongoing theory work and guide further refinement in these systems. In a productive partnership with the atomic theory group of Safronova et al. we have provided numerous new measurements in three-electron thallium and indium SafronovaMajumder2013 ; Ranjit2013 ; Augenbraun2016 ; Vilas2018 which have consistently tested the latest theory predictions at the 1% level or better. For the case of trivalent indium our precision was sufficient to test alternative theoretical methods based on a configuration interaction/all-orders approach (CI+all) Vilas2018 .
@@
This partnership has recently extended to the four-electron lead (Pb) system Porsev2016 ; Maser2019 . Lead, along with thallium and cesium Meekhof1993 ; Phipp1996 ; Vetter1995 ; Wood1997 , were the subject of experimental work in the 1990’s which all measured atomic parity nonconservation (PNC) at the 1% level of precision or better. The single-valence Cs system yields a combination of experimental and similarly precise atomic theory that has long provided an important low-energy test of fundamental electroweak theory Wood1997 ; Porsev2009 ; Dzuba2012 . At the time of the experimental work, the atomic theory of lead was a full order-of-magnitude less accurate than the experiment. Recent substantial progress in multi-valence theory work, however, has reduced the uncertainties in this system several fold Porsev2016 , so that even modest further improvements in theory accuracy will have significant impacts on the quality of future tests of fundamental physics in this arena.
@@
Another important class of symmetry-violation tests involving atoms and molecules has focused on searches for potential T-violating permanent electric dipole moments (EDMs) of the electron ChuppEDM ; Andreev2018 ; JILAEDM23 . Lead has played an important role as a partner in diatomic molecule-based searches using PbO and PbF DeMille2000 ; ShaferRay2006 ; Baklanov2010 . Very recently, it has been proposed Stuntz2024 that a diatomic molecule including lead as one partner (paired with a gold or silver atom) has excellent potential for a future cold-molecule-based EDM search. A first step to experimentally realizing such a cold diatomic system would be to produce laser-cooled samples of lead atoms. In parallel with the lead vapor cell Faraday spectroscopy work that is the subject of this manuscript, we have been pursuing lead atomic beam-based spectroscopy to measure polarizabilities and Stark-induced amplitudes. As part of that work we intend to demonstrate, for the first time, transverse laser cooling in this atomic system, a clear first-step to any eventual cold molecule production involving this species. We are thus motivated for many reasons to pursue and expand our atomic structure measurements in this heavy, multi-valence system.
In 2019, we measured for the first time the matrix element of the very weak electric quadrupole (E2) transition within the ground-state manifold of lead and found excellent agreement, at the 1% level of accuracy with the latest atomic theory work Maser2019 . This was among the first atomic structure benchmark measurements in this system with this level of precision. In the current manuscript we describe two new measurements of lead E1 transition amplitudes, both achieving better than 1% fractional accuracy, an improvement in precision of nearly an order of magnitude over previous experimental work in the lead system AlonsoMedina2001 , and we compare these results to the latest ab initio theory predictions from our theory partners, Porsev and Safronova Maser2019 ; SafronovaPC . Using a high-precision polarimetry technique similar to that used in Maser2019 , we have compared the transition strength of two electric dipole transitions originating from low-lying excited states to that of a ground-state magnetic dipole transition. The very small thermal Boltzmann population for the relevant excited-state levels reduces the observed E1 absorptivity of these strongly-allowed transitions so that they are comparable to the intrinsically weaker M1 transition originating from the ground state. In this manuscript, having outlined the atomic structure details in Sec. II, we describe the key apparatus features and summarize data collection and analysis in Secs III. Sec. IV discusses various potential systematic error contributions, and we conclude with a discussion of the results with a comparison to previous work and the latest theory as well as noting ongoing and future work.
@@ II Atomic line shapes and
The relevant energy levels and transitions are shown in Fig. 1. In this study, spectral features of the 368 nm and 406 nm E1 transitions are analyzed and compared with the 1279 nm M1 transition. While the E1 transitions are intrinsically much stronger than the forbidden M1 transition, at typical experimental temperatures ($\,{}^{\circ}$), the spectral features turn out to be quite comparable in size owing to the much smaller thermal populations of the ${}^{3}P_{1}$ and ${}^{3}P_{2}$ levels compared with ${}^{3}P_{0}$ (see Fig. 1). The majority of the experimental work was conducted by Faraday rotation spectroscopy, where our polarimetry system can detect milliradian optical rotations with microradian accuracy. At high sample temperatures, where there was well-resolved direct absorption, we used analysis of transmission spectra to complement the Faraday rotation analysis. Both are described here. We note that the M1 transition amplitude does not depend on detailed wave function information, and therefore is known very precisely Porsev2016 even for the case of this multi-valence atom. As such, it offers an ideal normalization tool in the experimental determination of these E1 amplitudes.
@@ II.1 Transmission analysis
where the homogeneous (FWHM) and Doppler ($1/e$) spectral widths are denoted by $\Gamma$ and $\sigma$ respectively. The resonant peak value of Eq. 2 will be a quantity that depends on the component line widths, and we refer to this as $\mathcal{V}(0)$. To be consistent with our definitions in Eq. 2, we then incorporate this amplitude factor into our definition of the peak optical depth. If we now consider the ratio of peak optical depths for our two transitions, various common factors cancel, including the (identical) path length through the vapor, and we can express the ratio of peak optical depths in our experiment as
@@
While the atomic density is the same for both observed spectra, the population ratio is not, and is given by the Boltzmann thermal population for each sub-level of the particular excited state: $N_{E1}/N_{M1}=e^{-{\Delta E}/k_{B}T}$, where $\Delta E$ is the relevant energy difference (for the case of the two relevant excited states in our current work, $\Delta E$ is 0.9695 eV for the ${{}^{3}}P_{1}$ state and 1.3205 eV for the ${{}^{3}}P_{2}$ state). From this expression it is immediately clear that accurate vapor cell temperature determination is a requisite for a precise transition amplitude determination.
The matrix element expressions suggested in Eq. 3 must be made explicit for each transition that we study, and, further, need to be connected to reduced matrix elements in atomic units to be compared to theoretical predictions. In our polarimetry apparatus atoms interact with linearly polarized light in all cases. This interaction is best viewed in a spherical basis anticipating the application of magnetic fields and Zeeman-shifted resonances. Appendix A makes this connection explicit for each of the three transitions we study, employing the Wigner-Eckart theorem and a related sum rule. For the simpler 368 nm E1 transition, as well as for the 1279 nm M1 transition where there is only one transition for a given circular polarization component, it is easy to show (see Appendix A) that the ratio of reduced matrix elements is equal to the ratio of the $\Delta m=1$ matrix elements. Thus we can find an expression that allows us to extract the E1 reduced matrix element from various experimentally measured and fitted parameters, as well as known constants:
@@
We have explicitly included the fundamental constants that put the matrix elements into atomic units: $e$ is the elementary charge, $a_{0}$ is the Bohr radius, $\mu_{B}$ is the Bohr magneton, and $c$ is the speed of light. The exponential factor and the final units conversion factor in Eq. 4 nearly (and fortuitously) cancel, so that the observed optical depths, as well as the Faraday rotation amplitudes discussed next, are quite comparable between the M1 and E1 transitions in all of our comparison experiments.
@@ II.2 Faraday rotation analysis
We finally turn to the more complicated 406 nm transition Faraday line shape. Referring again to the discussion in Appendix A as well as Fig. 2, this ($J=2\rightarrow J=1$) transition comprises a trio of $\sigma_{+}$ and corresponding $\sigma_{-}$ components each weighted according to the relevant Clebsch-Gordan coefficient and each Zeeman-shifted by different amounts according to the two different $g$-factors in play (labeled $g^{\prime}$ for the lower state and $g^{\prime\prime}$ for the upper state in Fig. 2). Thus we must expand our two-dispersion fitting function (Eq. 7) to include six terms with individual weighting factors and $g$-factors. Referring to the labels a, b, and c in Fig. 2, these weighting factors (which sum to unity) are: $w_{a}=6/10,w_{b}=3/10,w_{c}=1/10$ and the corresponding ‘effective’ $g$-factors are: $g_{a}=(2g^{\prime}-g^{\prime\prime}),\,g_{b}=g^{\prime},\,g_{c}=g^{\prime\prime}$. For analysis of the 406 nm transition Faraday data, we use this expanded, properly normalized six-dispersion curve model, and retaining the overall model amplitude factor $\mathcal{C}$ in our fitting process, we are then able to employ Eq. 8 as written to extract $\langle||E1||\rangle_{406}$.
Given that all of the Zeeman shifts here remain small compared to the overall dispersion curve linewidth, one can make an excellent two-dispersion approximate model for this transition, using the weighted average of each set of three dispersion curves and a single ‘weighted-average’ $g$-factor. This allows us to employ the same line shape model as for the 1279 and 368 transitions, and we find that the final results obtained using the ‘exact’ vs ‘approximate’ fits to the 406 transition Faraday data differ by less than 0.1% in all cases.
@@ III.1 Apparatus
The basic experimental apparatus is shown in Fig. 3 and is quite similar to that used in Maser2019 . In an effort to ensure temperature uniformity and stability for these new measurements, we placed a 15 cm-long quartz cell containing isotopically enriched 208Pb (purity$=99.97\%$) in the center of a home-built vacuum furnace. The cell windows are angled to avoid the possibility of multi-passing of the laser beams due to reflections. The furnace consists of three pairs of clamshell heaters surrounding a 1 m-long ceramic tube, evacuated and back-filled with 20 mTorr of argon. Each individual clamshell heater is driven by the amplified output of a function generator delivering a 10 kHz sine wave. The maximum total power output of the furnace is roughly 900 W. The large thermal mass of the furnace, in conjunction with a software p-i-d control loop which controls the amplitude of the function generator driving the heater amplifiers, ensures that after a warm-up period of several hours, the furnace temperature remains stable to better than $1\,^{\circ}$. The cell environment is monitored by 3 s-type (Pt:PtRh) precision thermocouples fixed at points A, B, and C within the stainless steel cylinder that encloses the Pb vapor cell (see Fig. 3). The thermocouple voltage is converted to a temperature reading using a Stanford Research Systems SRS 630 precision thermocouple reader, which corrects for dissimilar metal contact voltages at the instrument junctions by accurately monitoring the temperature of the connection pads. The 30 cm-long stainless cylinder and its endcaps help to provide a stable, uniform thermal environment for the vapor cell. Slight adjustments to the individual heater currents allow optimization of temperature uniformity. Thermocouple accuracy was confirmed (at least at lower temperatures) by putting them in contact with both a water/ice slurry and boiling water, which gave agreement with expected temperatures at the $0.1\,^{\circ}$ level. This is well within the quoted thermocouple accuracy over its full temperature range of ±$\pm{1.5}\,^{\circ}$. Surrounding the heaters are several layers of insulation, water cooling pipes, and a solenoid delivering 22.0(2) G/A. This was accurately determined (see Appendix B) by measuring the Zeeman splitting of the $(6p^{2})^{3}P_{1}$ level whose $g$-factor is accurately known Wood1968 . A $\mu$-metal cylinder surrounding the entire furnace limits any external transverse magnetic fields in the interaction region to less than 10 $\mu$T.
@@
The optical and signal processing setup is also similar to that of previous work in Pb Maser2019 , and is substantially the same for both the 368 nm and 406 nm E1 transition amplitude measurements. In each case, our 1279 nm laser (Sacher Lasertechnik) drives the reference M1 transition, while the beam from a home-built E1 external cavity diode laser is aligned using a dichroic mirror to traverse the same beam path as the 1279 nm laser. Path overlap was optimized by maximizing the signal enhancement from two-step excitation of the 1279 nm and 368 nm transition. (For the 406 nm transition, this enhancement was observed and optimized through a secondary effect, whereby the 1279 nm M1 laser enhances the ${}^{3}P_{1}$ level which further populates the ${}^{3}P_{2}$ level through collisions). We further ensure geometric overlap with the use of small collimating apertures placed 1.5 m apart through which both laser beams must pass. The lasers (which in the experiment are alternately blocked and unlocked using computer-controlled shutters) pass through the atomic sample located in between two Glan-Thompson polarizers. The emerging light beams are separated by a second dichroic mirror before being directed into their respective photodetectors. Between the first polarizer and the atomic sample is a magneto-optic element which consists of a high-Verdet-constant material. The principal difference in our two experimental setups was the use of two different crystals: a 1 cm-long, 1 cm-diameter TGG glass was used for the 406 nm measurement, and a 5 mm-long, 5 mm-diameter CeF3 unixial single crystal element for the 368 nm measurement. The latter was used for its favorable transmission properties in the UV Lacy2024 , though its twenty-fold smaller Verdet constant at IR wavelengths led to somewhat higher optical rotation noise levels for the M1 Faraday data here (yet still below $10\,\mu\mathrm{Rad}/\sqrt{\mathrm{Hz}}$). In each case, the relevant Faraday modulator crystal is subjected to a sinusoidally-varying (500 Hz) magnetic field of roughly 75 Gauss that modulates the light polarization. The detected signals are then demodulated using four lock-in amplifiers (Stanford Research Instruments, model SR810 DSP) referenced to $5001f=$ 500 Hz and $10002f=$ 1000 Hz for each laser. As discussed in Maser2019 ; Lacy2024 , the $2f$ component provides a signal directly proportional to the transmission signal (with no optical rotation information), while the ratio of the $1f$ and $2f$ lock-in signals gives a signal proportional to the atomic Faraday optical rotation, $\Phi_{F}(\omega)$ when we apply a longitudinal DC magnetic field to the atoms, so that:
@@
where $\kappa$ is a calibration constant determined experimentally in the course of data collection, as described below. Along with the $1f$ and $2f$ lock-in signals, a Fabry-Pérot transmission spectrum was collected to monitor the frequency scan of each laser. A 600 MHz resonant electro-optic modulator and RF synthesizer was used to determine the free spectral ranges (FSRs) of the M1 (500.7(3) MHz) and E1 (500.4(4) MHz) confocal cavities (not shown in Fig. 3).

### v2

@@ Abstract
We have completed measurements of two low-lying excited-state electric-dipole (E1) transition amplitudes in lead. Our measured reduced matrix elements of the $(6s^{2}6p^{2})^{3}P_{1}\to(6s^{2}6p7s)^{3}P_{0}$ transition at 368.3 nm and the 405.8 nm $(6s^{2}6p^{2})^{3}P_{2}\to(6s^{2}6p7s)^{3}P_{1}$ transition are 1.90(1) a.u. and 3.01(2) a.u. respectively, both measured to sub-1% precision and both in excellent agreement with the latest ab initio lead wavefunction calculations. These measurements were completed by comparing the low-field Faraday optical rotation spectra of each E1 transition in turn with that of the ground state ${}^{3}P_{0}-^{3}P_{1}$ M1 transition under identical experimental conditions. Our spectroscopy technique involves polarization modulation and lock-in detection yielding microradian-level optical rotation resolution. At temperatures where direct absorption was significant for both E1 and M1 transitions, we also extracted matrix element values from a direct optical absorption depth comparison. As part of this work we designed an interaction region within our furnace which allows precise measurement of our quartz vapor cell sample temperature and thus accurate determination of the Boltzmann thermal population of the low-lying excited states that were studied.
@@ I Introduction
Atoms and associated precision spectroscopic techniques have long been an excellent means to provide important, low-energy tests of fundamental physics associated with the standard model, as well as tests of potential new (“BSM”) physics beyond it DeMille2017 ; Safronova2018 . Heavy (high-$Z$, often multi-valence) atomic species in particular, due to their enhanced sensitivity to many of these phenomena, have played an important role in experimental searches. This high-precision experimental work has developed alongside improved ab initio atomic theory to link experimental results to interpretations and limits of the fundamental physics phenomena being targeted. Recent atomic theory work particularly targeted at multi-valence atomic systems, and broadly applicable to many such systems, has made notable progress in accuracy in recent years Safronova2009 ; Safronova2013 ; Safronova2025 . Unlike, for example, the alkali atom systems, far fewer general high-precision atomic structure measurements have been completed in heavy, multi-valence atoms with which to benchmark ongoing theory work and guide further refinement in these systems.
@@
Building on from previous work in trivalent systemsSafronovaMajumder2013 ; Ranjit2013 ; Augenbraun2016 ; Vilas2018 , our collaboration with the theory group of Safronova et al. has been recently extended to the four-electron lead (Pb) system Porsev2016 ; Maser2019 . Lead, along with thallium and cesium Meekhof1993 ; Phipp1996 ; Vetter1995 ; Wood1997 , were the subject of experimental work in the 1990’s which all measured atomic parity nonconservation (PNC) at the 1% level of precision or better. The single-valence Cs system yields a combination of experimental and similarly precise atomic theory that has long provided an important low-energy test of fundamental electroweak theory Wood1997 ; Porsev2009 ; Dzuba2012 . Recently, substantial progress in multi-valence theory work (including lead) has reduced the uncertainties in these systems several fold Porsev2016 , so that even modest further improvements in theory accuracy will have significant impacts on the quality of future tests of fundamental physics in this arena.
@@
Another important class of symmetry-violation tests involving atoms and molecules has focused on searches for potential T-violating permanent electric dipole moments (EDMs) of the electron ChuppEDM ; Andreev2018 ; JILAEDM23 . Lead has played an important role as a partner in diatomic molecule-based searches using PbO and PbF DeMille2000 ; ShaferRay2006 ; Baklanov2010 . Very recently, it has been proposed Stuntz2024 that a diatomic molecule including lead as one partner (paired with a gold or silver atom) has excellent potential for a future cold-molecule-based EDM search. A first step to experimentally realizing such a cold diatomic system would be to produce laser-cooled samples of lead atoms. In parallel with our lead vapor cell spectroscopy work, we intend to demonstrate, for the first time, transverse laser cooling in a lead atomic beam, the same apparatus which we will explore Stark-induced amplitudes and atomic polarizability measurements for additional high-precision tests of atomic theory. We are thus motivated for many reasons to pursue and expand our atomic structure measurements in this heavy, multi-valence system.
In 2019, we measured, for the first time, the matrix element of the very weak electric quadrupole (E2) transition within the ground-state manifold of lead and found agreement with the ab initio theory at the 1% level of accuracy Maser2019 . In the current manuscript, we describe two new measurements of lead E1 transition amplitudes, both of which with a fractional accuracy under 1%. These measurements offer an improvement in precision of nearly an order of magnitude over previous experimental work in the lead system AlonsoMedina2001 , and show excellent agreement with latest values from our theory partners Maser2019 ; SafronovaPC .
In this manuscript, having outlined the atomic structure details in Sec. II, we describe the key apparatus features and summarize data collection and analysis in Secs III. Sec. IV discusses various potential systematic error contributions, and we conclude with a discussion of the results with a comparison to previous work and the latest theory as well as noting ongoing and future work.
@@ II Atomic line shapes and
The relevant energy levels and transitions are shown in Fig. 1. In this study, spectral features of the 368 nm and 406 nm E1 transitions are analyzed and compared with the 1279 nm M1 transition. While the E1 transitions are intrinsically much stronger than the forbidden M1 transition, at typical experimental temperatures ($\,{}^{\circ}$), the spectral features turn out to be quite comparable in size owing to the much smaller thermal populations of the ${}^{3}P_{1}$ and ${}^{3}P_{2}$ levels compared with ${}^{3}P_{0}$ (see Fig. 1).
The majority of the experimental work was conducted by Faraday rotation spectroscopy, where our polarimetry system can detect milliradian optical rotations with microradian accuracy. At high sample temperatures, where there was well-resolved direct absorption, we used analysis of transmission spectra to complement the Faraday rotation analysis. Both are described here. We note that the M1 transition amplitude does not depend on detailed wave function information, and therefore is known very precisely Porsev2016 even for the case of this multi-valence atom. Indeed, using similar theoretical methods, M1 amplitude calculations by this theory group in other multi-valence systems has been shown to agree with experimental values at the 0.1% level of precision or better Brewer2019 ; Bothwell2025 . As such, the lead 1279 nm M1 transition offers an ideal normalization tool in our experimental determination of these E1 amplitudes.
@@ II.1 Transmission analysis
where the homogeneous (Lorentzian) and Doppler spectral widths are denoted by $\Gamma$ and $\sigma$ respectively. The resonant peak value of Eq. 2 will be a quantity that depends on the component line widths, and we refer to this as $\mathcal{V}(0)$. To be consistent with our definitions in Eq. 1, we then incorporate this amplitude factor into our definition of the peak optical depth. If we now consider the ratio of peak optical depths for our two transitions, various common factors cancel, including the (identical) path length through the vapor, and we can express this ratio as
@@
While the atomic density is the same for both observed spectra, the population ratio is not, and is given by the Boltzmann thermal population for each sub-level of the particular excited state: $N_{E1}/N_{M1}=e^{-{\Delta E}/k_{B}T}$, where $\Delta E$ is the relevant energy difference (for the case of the two relevant excited states in our current work, $\Delta E$ is 0.9695 eV for the ${{}^{3}}P_{1}$ state and 1.3205 eV for the ${{}^{3}}P_{2}$ state). From this expression, it is immediately clear that an accurate transition amplitude measurement will depend on precise cell temperature determination.
The matrix element expressions suggested in Eq. 3 must be made explicit for each transition that we study, and, further, need to be connected to reduced matrix elements in atomic units to be compared to theoretical predictions. In our polarimetry apparatus, atoms interact with linearly polarized light in all cases. This interaction is best viewed in a spherical basis anticipating the application of magnetic fields and Zeeman-shifted resonances. Appendix A makes this connection explicit for each of the three transitions we study, employing the Wigner-Eckart theorem and a related sum rule. For the simpler 368 nm E1 transition, as well as for the 1279 nm M1 transition where there is only one transition for a given circular polarization component, it is easy to show (see Appendix A) that the ratio of reduced matrix elements is equal to the ratio of the $\Delta m=1$ matrix elements. Thus we can find an expression that allows us to extract the E1 reduced matrix element from various experimentally measured and fitted parameters, as well as known constants:
@@
We have explicitly included the physical constants that put the matrix elements into atomic units: $e$ is the elementary charge, $a_{0}$ is the Bohr radius, $\mu_{B}$ is the Bohr magneton, and $c$ is the speed of light. The exponential factor and the final units conversion factor in Eq. 4 nearly (and fortuitously) cancel, so that the observed optical depths, as well as the Faraday rotation amplitudes discussed next, are quite comparable between the M1 and E1 transitions in all of our comparison experiments.
@@ II.2 Faraday rotation analysis
We finally turn to the more complicated 406 nm transition Faraday line shape. Referring again to the discussion in Appendix A as well as Fig. 2, this ($J_{i}=2\rightarrow J_{f}=1$) transition comprises a trio of $\sigma_{+}$ and corresponding $\sigma_{-}$ components each weighted according to the relevant Clebsch-Gordan coefficient and each Zeeman-shifted by different amounts according to the two different $g$-factors in play (labeled $g^{\prime}$ for the lower state and $g^{\prime\prime}$ for the upper state in Fig. 2). Thus we must expand our two-dispersion fitting function (Eq. 7) to include six terms with individual weighting factors and $g$-factors. Referring to the labels a, b, and c in Fig. 2, these weighting factors (which sum to unity) are: $w_{a}=6/10,w_{b}=3/10,w_{c}=1/10$ and the corresponding ‘effective’ $g$-factors are: $g_{a}=(2g^{\prime}-g^{\prime\prime}),\,g_{b}=g^{\prime},\,g_{c}=g^{\prime\prime}$. For analysis of the 406 nm transition Faraday data, we use this expanded, properly normalized six-dispersion curve model, and retaining the overall model amplitude factor $\mathcal{C}$ in our fitting process, we are then able to employ Eq. 8 as written to extract $\langle||E1||\rangle_{406}$.
Given that all of the Zeeman shifts here remain small compared to the overall dispersion curve linewidth, one can make an excellent two-dispersion approximate model for this transition, using the weighted average of each set of three dispersion curves and a single ‘weighted-average’ $g$-factor. This allows us to employ the same line shape model as for the 1279 nm and 368 nm transitions, and we find that the final results obtained using the ‘exact’ vs ‘approximate’ fits to the 406 nm transition Faraday data differ by less than 0.1% in all cases.
@@ III.1 Apparatus
The basic experimental apparatus is shown in Fig. 3 and is quite similar to that used in Maser2019 . In an effort to ensure temperature uniformity, we placed a 15 cm-long quartz cell containing isotopically enriched 208Pb (purity$=99.97\%$) in the center of a home-built vacuum furnace. The cell windows are angled to avoid the possibility of multi-passing of the laser beams due to reflections. The furnace consists of three pairs of clamshell heaters surrounding a 1 m-long ceramic tube, evacuated and back-filled with 20 Torr of argon. Each individual clamshell heater is driven by the amplified output of a function generator delivering a 10 kHz sine wave. The maximum total power output of the furnace is roughly 900 W. The large thermal mass of the furnace, in conjunction with a software p-i-d control loop which controls the amplitude of the function generator driving the heater amplifiers, ensures that after a warm-up period of several hours, the furnace temperature remains stable to better than $1\,^{\circ}$.
The cell environment is monitored by 3 s-type (Pt:PtRh) precision thermocouples fixed at points A, B, and C within the stainless steel cylinder that encloses the Pb vapor cell (see Fig. 3). The thermocouple voltage is converted to a temperature reading using a Stanford Research Systems SRS 630 precision thermocouple reader, which corrects for dissimilar metal contact voltages at the instrument junctions by accurately monitoring the temperature of the connection pads. The 30 cm-long stainless cylinder and its endcaps help to provide a stable, uniform thermal environment for the vapor cell. Slight adjustments to the individual heater currents allow optimization of temperature uniformity. Thermocouple accuracy was confirmed (at least at lower temperatures) by putting them in contact with both a water/ice slurry and boiling water, which gave agreement with expected temperatures at the $0.1\,^{\circ}$ level. This is well within the quoted thermocouple accuracy over its full temperature range of ±$\pm{1.5}\,^{\circ}$. Thermocouple variance was tested at a range of temperatures by clamping them to a single point, but no difference in thermocouple reading was observed beyond the ±$\pm{0.5}\,^{\circ}$ level at any time during the heating cycle.
Surrounding the heaters are several layers of insulation, water cooling pipes, and a solenoid delivering 22.0(2) G/A. This was accurately determined (see Appendix B) by measuring the Zeeman splitting of the $(6p^{2})^{3}P_{1}$ level whose $g$-factor is known to high precision Wood1968 . A $\mu$-metal cylinder surrounding the entire furnace limits any external transverse magnetic fields in the interaction region to less than 10 $\mu$T.
The optical and signal processing setup is also similar to that of previous work in Pb Maser2019 , and is substantially the same for both the 368 nm and 406 nm E1 transition amplitude measurements. In each case, our 1279 nm laser (Sacher Lasertechnik) drives the reference M1 transition, while the beam from a home-built E1 external cavity diode laser is aligned using a dichroic mirror to traverse the same beam path as the 1279 nm laser. Path overlap was optimized by maximizing the signal enhancement from two-step excitation of the 1279 nm and 368 nm transition. (For the 406 nm transition, this enhancement was observed and optimized through a secondary effect, whereby the 1279 nm M1 laser enhances the ${}^{3}P_{1}$ level which further populates the ${}^{3}P_{2}$ level through collisions). We further ensure geometric overlap with the use of small collimating apertures placed 1.5 m apart through which both laser beams must pass.
@@
The lasers (which in the experiment are alternately blocked and unblocked using computer-controlled shutters) pass through the atomic sample located in between two Glan-Thompson polarizers. The emerging light beams are separated by a second dichroic mirror before being directed into their respective photodetectors. Between the first polarizer and the atomic sample is a magneto-optic element which consists of a high-Verdet-constant material. The principal difference in our two experimental setups was the use of two different crystals: a 1 cm-long, 1 cm-diameter TGG glass was used for the 406 nm measurement, and a 5 mm-long, 5 mm-diameter CeF3 unixial single crystal element for the 368 nm measurement. The latter was used for its favorable transmission properties in the UV Lacy2024 , though its twenty-fold smaller Verdet constant at IR wavelengths led to somewhat higher optical rotation noise levels for the M1 Faraday data here (yet still below $10\,\mu\mathrm{Rad}/\sqrt{\mathrm{Hz}}$).
In each case, the relevant Faraday modulator crystal is subjected to a sinusoidally-varying (500 Hz) magnetic field of roughly 75 Gauss that modulates the light polarization. The detected signals are then demodulated using four lock-in amplifiers (Stanford Research Instruments, model SR810 DSP) referenced to $5001f=$ 500 Hz and $10002f=$ 1000 Hz for each laser. As discussed in Maser2019 ; Lacy2024 , the $2f$ component provides a signal directly proportional to the transmission signal (with no optical rotation information), while the ratio of the $1f$ and $2f$ lock-in signals gives a signal proportional to the atomic Faraday optical rotation, $\Phi_{F}(\omega)$ when we apply a longitudinal DC magnetic field to the atoms, so that:
@@
where $\kappa$ is a calibration constant determined experimentally in the course of data collection, as described below.
Along with the $1f$ and $2f$ lock-in signals, a Fabry-Pérot transmission spectrum was recorded to provide the relative frequency axis of each collected Faraday spectrum. A 600 MHz resonant electro-optic modulator and RF synthesizer was used to determine the free spectral ranges (FSRs) of the M1 (500.7(3) MHz) and E1 (500.4(4) MHz) confocal cavities (not shown in Fig. 3).